OpenStudy (anonymous):

find a sub 15 for the sequence -1.5,3, -6...

OpenStudy (anonymous):

The sequence is as follows:\[-\frac{ 3 }{ 2 }, 3, -6, ....\] What pattern does the sequence follow?

OpenStudy (anonymous):

Geometric?

OpenStudy (anonymous):

\[-\frac{ 3 }{ 2 },\frac{ 6 }{ 2 },-\frac{ 12 }{ 2 }, \frac{ 24 }{ 2 }, -\frac{ 48 }{ 2 },...\]The sequence follows a linear equation. See how you're just multiplying each term by -2? You can figure out an equation for the whole series.

OpenStudy (anonymous):

Woah. Mind = Blown!

OpenStudy (anonymous):

\[y=-2x\] Is that it?

OpenStudy (anonymous):

It's usually described as something along the lines of \[a_{n}=-\frac{3}{2}(-2)^{n}\]

OpenStudy (anonymous):

I'm sorry, it might be a little advanced for your level, you probably wouldn't need to find the equation, but it goes something like this: \[a_n = 3 (-1)^n 2^{n-2}\] I think for your level, your teacher probably just wanted you to count to the 15th term. if \[a _{1}=-1.5\] \[a _{2}=3\] \[a _{3}=-6\] count up to \[a _{15}=?\] and check it with the equation I gave you, plugging in n=15.

OpenStudy (anonymous):

So I should come to the answer of 49152 correct?

OpenStudy (anonymous):

I think you went 1 too many. 1) -1.5 2) 3 3) -6 4) 12 5) -24 6) 48 7) -96 8) 192 9) -384 10) 768 11) -1536 12) 3072 13) -6144 14) 12288 15) -24576 16) 49152

OpenStudy (anonymous):

\[a_{15}=3(−1)^{15}2^{15−2}=-24576\] Tell your teacher you solved for that equation in your head, maybe they'll graduate you early, haha

OpenStudy (anonymous):

Haha! I wish!

OpenStudy (agent0smith):

@chrismoon looks like \[a _{n } = \frac{ 3 }{ 4 }(-2)^n \] should work too. Though the standard form for geometric (which i was trying to remember when i worked that out) is \[a _{n} = a _{1} r ^{n-1}\] where a1 is the first term and r is the common ratio, which would be: \[a _{n} = -1.5 (-2) ^{n-1}\] which works too.

OpenStudy (agent0smith):

@SOSAlgebra6629 have you done the standard form for geometric sequence? This form: \[a _{n} = a _{1} r ^{n-1}\]

OpenStudy (anonymous):

No. What is that?

OpenStudy (anonymous):

Yea, they're the same equation, I just didn't simplify. Since\[2^{n-2}=\frac{ 2^n }{ 2^2 }\]and \[(-1)^{n}2^n=(-2)^n\] My original equation can be reduced to yours.

OpenStudy (anonymous):

Wait a second, is the a sub 1 the first number in the sequence? Is the r the rate of change?

OpenStudy (agent0smith):

^yes to both. r is the common ratio which is -2.

OpenStudy (agent0smith):

if you've done geometric sequences you've prob used it\[\large a _{n} = a _{1} r ^{n-1}\]a sub n just means the n-th term.